OpenStudy (anonymous):

find the gcf 12, 18, and 84 can someone helpme with this?

OpenStudy (mathstudent55):

To find a GCF, you must first find the prime factors of all the numbers. Do you know how to do that?

OpenStudy (anonymous):

no im not sure

OpenStudy (mathstudent55):

Ok, then this is how it's done. A prime number is a number greater than one that is divisible by only itself and 1. The first several prime numbers are: 2, 3, 5, 7, 11, 13, 17, 19, 23, 29. To find the prime factors of a number, try to divide the number by the smallest prime you can divide it by. Do it as many times as you can. Then move on to the next prime number, and so on, until you get the division to equal 1. L:et's start with 12. Divide it by 2: 12/2 =6, now divide 6 by 2, 6/2 = 3. 3 can't be divided by 2, so divide it by 3: 3/3 = 1 That means the prime factors of 12 are 2,2, and 3. Check: 2 x 2 x 3 = 12 Now try that with 18 and 84

OpenStudy (mathstudent55):

One way of doing it is with a "factoring tree." See 12 done below.